# Mind: Towards Immersive Psychological Healing with Multi-agent Inner Dialogue

### v1

@@ 3.1 Setup
3.2 Prelinimary: SP Role-playing Evaluation
The results are shown in Table 2 and the evaluation metrics is shown in Appendix B. The Gemini-2.0-flash model emerged as the best performer among the tested LLMs. In comparison, GPT-4o performed well in some areas but lagged in Emotional Expression and Personalization. Models like GPT-3.5-Turbo, Llama-3.1-8B-Instruct, and Deepseek-R1 showed weaker results, particularly in emotional and personalized responses. Qwen2.5 models performed the poorest, especially in emotional expression and accuracy, with scores not surpassing 3.2 in any dimension.
@@ 3.2 Prelinimary: SP Role-playing Evaluation
3.3 Main Results
The mean scores of each paradigm are shown in Figure 4. Mind demonstrated significant strengths in all six core assessment dimensions. Quantitative analysis showed that our paradigm performed particularly well on the dimensions of interest and satisfaction, reaching a perfect score of 5, compared to all the baseline methods of traditional counseling, traditional empathy training, and chatbot. Notably, in terms of the engagement index, Mind achieved an absolute improvement of 17.1% over the suboptimal method of traditional counseling, which reflects the increased motivation of the caller users that Mind can improve, so that they cooperate and participate in psychotherapy. On the dimensions of immersion, coherence and emotional relief, MIND also outperforms/equals the remaining three paradigms, which fully demonstrates that Mind has the potential to advance psychological interventions by combining the scalability of LLMs with human-centered interaction design.
4 Analysis
4.1 Thematic Scenarios Ablation
This framework is applicable to a variety of thematic scenarios, including but not limited to work, family, and interpersonal issues. To analyze the differences in effectiveness across different themes within this framework, we independently generated five examples for each of the seven themes in the C2D2 dataset. Similarly, we invited evaluators with psychological therapy expertise to score these examples. As shown in Table 3, the performance of different themes varies under our framework. Most themes perform well in “Immersion” and “Coherence,” indicating that the system effectively engages users and maintains logical consistency. Emotional relief and satisfaction are high, especially in themes like “Work issues” and “Discrepancy between ideal and reality.” However, “Random negative events” and “Family issues” score lower in certain dimensions, such as engagement and interest, and may require further optimization.
4.2 Agent Involvement Ablation
Our framework consists of four agents: trigger, devil, guide, and strategist. To evaluate the effectiveness of Mind’s two core agents (i.e., the guide and strategist) as well as the memorization mechanism, we conducted several ablation experiments to assess their impact on user experience and demonstrate the importance of each component. Specifically, we randomly generated three examples for each ablation experiment. We recruited 7 clinical psychology researchers with professional expertise to evaluate six content evaluation metrics, as outlined in Table 1. The experimental results, presented in Figure 5, show that each agent significantly contributes to the overall framework. The removal of any agent or the memorization mechanism notably diminishes the quality of the generated content, underscoring the collective importance of all agents in the framework.
4.3 Case Study
To demonstrate Mind’s effectiveness in real-world applications, we present a case study in Appendix D, featuring a four-round dialogue on the theme of “work issues,” with the concern: “Despite studying hard, my grades remain poor, and effort seems useless in a talent-driven society.” The case study shows how the devil agent gains confidence through the player’s comforting words, while the player also develops greater self-compassion and reconciles with their own concerns.
5 Related Work
5.1 LLM Agent
5.2 LLM-assisted Psychology
6 Conclusion
Ethics Statement
The system used in this study is not intended to replace professional psychological treatment but rather to provide an effective option for clinical therapy. Before deployment, it is essential to ensure the presence of licensed professionals for supervision. Our evaluation method ensures the participation of mental health professionals aged 18 and above. The human evaluators’ ages range from 25 to 45 years, and their professions include one psychiatrist, two rehabilitation therapists, two psychotherapists, and two nurses. Prior to the experiment, we provided the human evaluators with detailed experimental guidelines.
Limitations
This framework has been evaluated exclusively in a Chinese-language context, which poses a limitation in terms of localizing psychological healing applications for different linguistic and cultural settings. While this study represents a significant step forward in shifting the paradigm of psychological healing, moving beyond the focus on training LLMs specifically for the psychological domain., it remains an initial attempt. To effectively implement this research into everyday psychological therapy, further extensive studies and clinical trials involving real mental health patients are necessary. Additionally, the framework’s guide agent could benefit from being replaced with a more specialized therapeutic model, which could enhance the system’s performance. Moreover, the framework used in this study is a simplified prototype. In the original theory , characters interact within a VR setting. There is significant potential for expanding this framework into more sophisticated formats, such as VR-based applications, to provide users with a more immersive and enriching therapeutic experience. Further exploration is required to address challenges related to the scalability of the system across various therapeutic scenarios and languages. Additionally, it remains unclear how the integration of this framework will scale in real-world settings with diverse patient populations, which presents another area for future research.
Appendix A Baseline Methods
This section provides a comprehensive overview of the baseline methods that we have employed. These methods serve as the foundational approaches in our study, and we introduce two distinct LLM-based baselines: (1) Chat-Bot; (2) Traditional Empathy Training.
Chat-Bot employs a simulated psychologist agent to engage in communication with patients suffering from cognitive distortions. During the conversation, it identifies the types of cognitive distortions and provides comfort and cognitive restructuring to the patients.
Traditional Empathy Training employs role reversal in four phases to address cognitive distortions. In Phase 1, self-critical participants interact with a crying child avatar as an adult, demonstrating empathy. In Phase 2, some participants switch to the child avatar to receive comfort from their past selves, while others observe from a third-person perspective as a control. Phase 3 involves adapting to new perspectives: first-person participants embody the child avatar, while third-person participants observe without a virtual body. In Phase 4, participants re-experience empathy from the child’s perspective, with real-time replays of the adult’s gestures and voice.To better align with our current work, we simulated this process using LLMs. An agent, describing actions, demeanor, and emotions, played the role of the crying child. Participants provided verbal comfort and interacted with the agent, observing changes in the crying child. Once the interaction concluded (i.e., when the crying child stopped crying), the comforter assumed the child’s perspective to review their comforting words and the child’s responses, describing their psychological state. This approach, using agents, replicated the role reversal process typically conducted in Virtual Reality (VR), with specific prompts detailed in Appendix C.
Appendix B SP Role-playing Assessment
(1) Dialogue Stability
(2) Language Fluency
(3) Emotional Expression
(4) Personalization & Diversity
(5) Accuracy
Appendix C Prompt Templates

### v2

@@ 3.1 Setup
3.2 SP Role-playing Evaluation
We begin with a preliminary role-playing experiment to assess the performance of various models in the Simulated Patient (SP) role-playing task. The results are presented in Table 1. Among these models, Gemini-2.0-flash performed best overall. While GPT-4o showed strength in some areas, it fell short in Emotional Expression and Personalization. Models such as GPT-3.5-Turbo, Llama-3.1-8B-Instruct, and Deepseek-R1 delivered weaker performance, especially in emotional and personalized responses. Qwen2.5 models ranked lowest, scoring below 3.2 across all dimensions, particularly in emotional expression and accuracy. Based on these findings, we select Gemini-2.0-flash for our main experiments due to its superior handling of the role-playing task and overall robustness.
Theme | IM | CO | EN | ER | SA | IN |
Work issues | 3.25 | 3.00 | 3.50 | 3.25 | 3.75 | 3.75 |
Random negative events | 3.25 | 3.50 | 3.25 | 3.75 | 3.50 | 3.50 |
Interpersonal issues | 4.25 | 3.75 | 4.25 | 3.75 | 4.25 | 4.00 |
Economic issues | 3.00 | 4.00 | 3.25 | 3.75 | 3.75 | 3.00 |
Family issues | 4.00 | 3.75 | 3.50 | 3.75 | 3.50 | 3.75 |
Physical stress | 4.25 | 4.25 | 3.75 | 3.75 | 3.75 | 3.75 |
Discrepancy between ideal and reality | 4.25 | 4.00 | 3.75 | 3.75 | 4.50 | 3.00 |
Model | EmoLLM | CACTUS | MIND | Control Group | ||||
Positive | Negative | Positive | Negative | Positive | Negative | Positive | Negative | |
Average Fluctuation | 0.36 | -0.11 | 1.35 | -0.52 | 1.46 | -0.65 | -0.11 | 0.03 |
Metrics | EmoLLM | CACTUS | MIND |
IM | 2.5 | 3.5 | 5.0 |
CO | 2.5 | 4.5 | 4.5 |
EN | 2.0 | 4.0 | 4.5 |
ER | 2.5 | 3.5 | 5.0 |
SA | 2.0 | 4.0 | 5.0 |
IN | 2.0 | 3.5 | 4.5 |
3.3 Main Results
The mean scores of each paradigm are shown in Figure 4. Mind demonstrated significant strengths in all six core assessment dimensions. Quantitative analysis showed that our paradigm performed particularly well on the dimensions of interest and satisfaction, reaching a perfect score of 5, compared to all the baseline methods of traditional counseling, traditional empathy training, and chat-bot. Notably, in terms of the engagement index, Mind achieved an absolute improvement of 17.1% over the suboptimal method of traditional counseling, which reflects the increased motivation of the caller users that Mind can improve, so that they cooperate and participate in psychotherapy. On the dimensions of immersion, coherence and emotional relief, Mind also outperforms/equals the remaining three paradigms, which fully demonstrates that Mind has the potential to advance psychological interventions by combining the scalability of LLMs with human-centered interaction design.
While closed-source models such as Gemini-2.0-flash showed a slight advantage—likely attributable to their state-of-the-art capabilities in reasoning, cognition, and role simulation—the framework also enables open-source models to achieve effective therapeutic interactions, as is shown in Figure 4. We therefore emphasize that our methodology is not tied to any single model: it leverages the scalability of LLMs together with human-centered interaction design, making it broadly applicable across both closed-source and open-source systems for advancing psychological interventions.
3.4 Human Experiment
We recruited a total of 8 volunteers (3 males and 5 females) with similar age, educational background, and living conditions. To assess changes in clients’ emotional states, we employed the Positive and Negative Affect Schedule (PANAS) questionnaire (Watson et al., 1988), which comprises 20 items covering 10 dimensions of positive and 10 dimensions of negative affect. Detailed experimental settings and the full PANAS questionnaire are provided in Table 6.
We calculated the average positive and negative emotional fluctuations of participants when interacting with three different systems: EmoLLM (Yang et al., 2024), CACTUS (Lee et al., 2024), Mind, and a control group. The results are reported in Table 3. In addition, participants rated each system across six subjective evaluation dimensions, with scores summarized in Table 4.
As shown, Mind achieved the best overall performance, outperforming other systems both in terms of emotional improvement measured by PANAS and in subjective ratings across all six evaluation criteria.
4 Analysis
4.1 Thematic Scenarios Ablation
This framework is applicable to a variety of thematic scenarios, including but not limited to work, family, and interpersonal issues. To analyze the differences in effectiveness across different themes within this framework, we independently generated five examples for each of the seven themes in the C2D2 dataset. Similarly, we invited evaluators with psychological therapy expertise to score these examples. As shown in Table 2, the performance of different themes varies under our framework. Most themes perform well in “Emotional Relief” and “Satisfaction”, indicating that the system can significantly alleviate users’ emotions, fully exert its healing effects, and provide users with a positive experience. Immersion and Engagement are high, especially in themes like “Physical stress” and “Interpersonal issues”. However, “Work issues” and “Economic issues” score lower in certain dimensions, which may require further optimization.
4.2 Agent Involvement Ablation
Our framework consists of four agents: trigger, devil, guide, and strategist. To evaluate the effectiveness of Mind’s two core agents (i.e., the guide and strategist) as well as the memorization mechanism, we conducted several ablation experiments to assess their impact on user experience and demonstrate the importance of each component. Specifically, we randomly generated three examples for each ablation experiment. We recruited 4 clinical psychology researchers with professional expertise to evaluate six content evaluation metrics, as outlined in Table 9.
The experimental results are presented in Figure 5, which shows that each agent significantly contributes to the overall framework. The removal of any agent or the memorization mechanism notably diminishes the quality of the generated content, underscoring the collective importance of all agents in the framework. Specifically, relying solely on the guide or the devil agent also leads to a decline in framework effectiveness, further underscoring the advantages of multi-agent collaboration.
4.3 Strategist with Structured Facilitation Protocol
One potential concern is the absence of an explicit facilitation protocol to manage stagnant conversations, which could risk inconsistent therapeutic progress in real-world applications. To examine this issue, we incorporated a structured facilitation protocol into the strategist agent’s prompt design following (Seo et al., 2014), with details provided in Appendix F. We then compared outcomes generated with and without this protocol using Gemini-2.0-flash. Across 70 samples (10 per theme for 7 themes), the proportion of failed cases-defined as instances where user thoughts were not fully shifted within 10 rounds-decreased only slightly, from 9% to 6%. This minimal improvement suggests that while the structured facilitation protocol provides an additional safeguard against conversational stagnation, its overall influence on performance is relatively limited.
4.4 Case Study
We present a case study in Appendix G, featuring a four-round dialogue on the theme of “work issues,” with the concern: “Despite studying hard, my grades remain poor, and effort seems useless in a talent-driven society.” The case study shows how the devil agent gains confidence through the player’s comforting words, while the player also develops greater self-compassion and reconciles with their own concerns.
5 Related Work
5.1 LLM Agent
5.2 LLM-assisted Psychology
6 Conclusion
Ethics Statement
The system used in this study is not intended to replace professional psychological treatment but rather to provide an effective option for clinical therapy. Before deployment, it is essential to ensure the presence of licensed professionals for supervision. Our evaluation method ensures the participation of mental health professionals and human experiment participants aged 18 and above. The human evaluators’ ages range from 25 to 45 years, and their professions include one psychiatrist, two rehabilitation therapists, two psychotherapists, and two nurses. The human experiment participants’ ages range from 18 to 21 years, and are all university students. Prior to the experiment, we provided the human evaluators with detailed experimental guidelines.
Limitations
While this study represents a significant step forward in shifting the paradigm of psychological healing, moving beyond the focus on training LLMs specifically for the psychological domain., it remains an initial attempt. To effectively implement this research into everyday psychological therapy, further extensive studies and clinical trials involving real mental health patients are necessary. Additionally, the framework’s guide agent could benefit from being replaced with a more specialized therapeutic model, which could enhance the system’s performance. Moreover, the framework used in this study is a simplified prototype. In the original theory , characters interact within a VR setting. There is significant potential for expanding this framework into more sophisticated formats, such as VR-based applications, to provide users with a more immersive and enriching therapeutic experience. Further exploration is required to address challenges related to the scalability of the system across various therapeutic scenarios and languages. Additionally, it remains unclear how the integration of this framework will scale in real-world settings with diverse patient populations, which presents another area for future research.
Acknowledgments
We are appreciated to all patients and their legal guardians in the study. We thank Yiran Zhao (CTO, Shuyunxushi Technologies Co., Ltd., China) for research design and hardware support. We sincerely thank Xingmei Quan (Co-founder of Xinyan Community Psychological Service Development Center) for generously sharing her professional expertise in Cognitive Behavioral Therapy to support this study. We also acknowledge the platform support provided by the SCU Brain Health Research and Innovation Network research group (BRAINET).
Appendix A Baseline Methods
This section provides an overview of the baseline methods employed in our study. These methods serve as fundamental points of comparison, including two LLM-based baselines—Chat-Bot and Traditional Empathy Training—and one human-centered baseline, Traditional Counseling.
Chat-Bot refers to an LLM configured for psychological healing, simulating a virtual therapist that engages with users to provide emotional support. It identifies cognitive distortions during the conversation, offers comfort, and attempts cognitive restructuring. This baseline serves as a lower bound, reflecting the performance of current single-agent psychological healing LLMs without the structured multi-agent framework proposed in our work.
Traditional Empathy Training employs role reversal in four phases to address cognitive distortions. In Phase 1, self-critical participants interact with a crying child avatar as an adult, demonstrating empathy. In Phase 2, some participants switch to the child avatar to receive comfort from their past selves, while others observe from a third-person perspective as a control. Phase 3 involves adapting to new perspectives: first-person participants embody the child avatar, while third-person participants observe without a virtual body. In Phase 4, participants re-experience empathy from the child’s perspective, with real-time replays of the adult’s gestures and voice.To better align with our current work, we simulated this process using LLMs. An agent, describing actions, demeanor, and emotions, played the role of the crying child. Participants provided verbal comfort and interacted with the agent, observing changes in the crying child. Once the interaction concluded (i.e., when the crying child stopped crying), the comforter assumed the child’s perspective to review their comforting words and the child’s responses, describing their psychological state. This approach, using agents, replicated the role reversal process typically conducted in Virtual Reality (VR), with prompts detailed in Appendix F.
In our prompt design, the role of a little girl is adopted as the main character because this figure has been widely used in traditional VR-based empathy training. Moreover, it offers flexibility for personalization based on the patient’s needs-for instance, it can be adapted to a little boy, an adult woman, or an adult man. This adaptability helps reduce the patient’s psychological defensiveness while fostering empathy and a sense of care. As shown in Table 5, the variance in results across different character roles is relatively small, suggesting that the choice of role does not significantly affect the overall outcomes of the experiment.
Traditional Counseling refers to in-person or real-time professional psychotherapy, conducted by licensed human therapists. This condition serves as the gold standard baseline for human evaluation and is used for benchmarking the perceived effectiveness of our framework.
Appendix B SP Role-playing Assessment
(1) Dialogue Stability
(2) Coherence & Fluency
(3) Emotional Expression
(4) Personalization & Diversity
(5) Accuracy
Metric | Character | Average | Standard Deviation |
Immersion | little girl | 4.00 | |
little boy | 3.75 | ||
woman | 3.25 | 0.53 | |
man | 3.25 | ||
self in mirror image | 4.50 | ||
Coherence | little girl | 4.00 | |
little boy | 3.25 | ||
woman | 4.00 | 0.31 | |
man | 3.75 | ||
self in mirror image | 3.75 | ||
Engagement | little girl | 3.50 | |
little boy | 3.25 | ||
woman | 3.50 | 0.47 | |
man | 3.00 | ||
self in mirror image | 4.25 | ||
Emotional Relief | little girl | 3.75 | |
little boy | 3.50 | ||
woman | 3.50 | 0.45 | |
man | 3.00 | ||
self in mirror image | 4.25 | ||
Satisfaction | little girl | 4.00 | |
little boy | 3.50 | ||
woman | 3.25 | 0.54 | |
man | 3.25 | ||
self in mirror image | 4.50 | ||
Interest | little girl | 3.50 | |
little boy | 3.00 | ||
woman | 4.00 | 0.50 | |
man | 3.00 | ||
self in mirror image | 4.00 |
Appendix C Human Experiment Details
C.1 Pipeline
Step 1: Participant Recruitment and Screening
We recruited a total of 8 volunteers, 3 males and 5 females, and similar in age, educational background, and living situation. There were 2 volunteers who did not participate in the model interaction and were only recruited to compare the likelihood of natural fluctuations in mood over time. Participants were required to have worries that bothered them for 1 day to 1 week. We paid 50rmb per participant as a subsidy.
Step 2: Pre-test Evaluation
We measured clients’ mood changes using the Positive and Negative Affect Scale (PANAS) questionnaire. The questionnaire contains 20 questions covering 10 positive and 10 negative emotion dimensions. Prior to the start of the experiment, our coauthor mental health experts introduced the PANAS questionnaire and the scoring criteria of the six dimensions we proposed, and informed participants that they could terminate the experiment at any time, and that the experimental data would be kept completely confidential and anonymized, so that they should fill in the form as honestly and as naturally as possible, and give the feedback that most closely corresponded to their inner thoughts.
Step 3: Experimental Design and Model Assignment
We randomly assigned the six participants in the experimental group to three systems (EmoLLM, CACTUS, Mind), with each system corresponding to two clients. The control group did not engage in any dialog and only filled out the questionnaire twice (30 minutes apart).
Step 4: Experimental implementation
Participants entered a real counseling room and engaged in five rounds of text-based conversations with the assigned model via a computer. The content of the conversations was kept strictly confidential, and the model stopped recording as soon as the conversations were over. Participants were asked to communicate about their “short-term negative experiences”, such as academic stress, relationship problems, and so on.
Step 5: Post-test and data collection
After the dialogues, participants completed the PANAS questionnaire again to compare the change in mood (e.g., whether the negative mood score decreased). All the emotion ratings are presented in Table 7 And anonymous feedback on the system was collected. We collected the results as follows:
Client 5 stated that MIND is helpful for emotion channeling, and that it can give a positive suggestion to oneself by consoling others. Client 3 affirmed CACTUS’s emotion channeling ability, but said that the response style is “a little bit formatted and not very flexible”, and Client 1 is skeptical about EmoLLM, thinking that “the content is empty, a lot of words, but there is no useful information, with low emotional value, unattractive”. Client 2 and Client 6 mentions the “problem of long response time”, which may be a major constraint to the application of LLM in real-life counseling scenarios.
Positive and Negative Affect Schedule (PANAS) | |
I. Positive Affect | |
1. | Interested |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
2. | Excited |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
3. | Strong |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
4. | Enthusiastic |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
5. | Proud |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
6. | Alert |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
7. | Inspired |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
8. | Determined |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
9. | Attentive |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
10. | Active |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
II. Negative Affect | |
11. | Distressed |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
12. | Upset |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
13. | Guilty |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
14. | Scared |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
15. | Hostile |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
16. | Irritable |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
17. | Ashamed |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
18. | Nervous |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
19. | Jittery |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much | |
20. | Afraid |
A. Very Rarely or Not at All B. Very Little C. Moderately D. Quite a Bit E. Very Much |
Emotion | EmoLLM | CACTUS | MIND | Control Group | ||||||||||||||||||||
client1 | client2 | client3 | client4 | client5 | client6 | client7 | client8 | |||||||||||||||||
b | a | b | a | b | a | b | a | b | a | b | a | b | a | b | a | |||||||||
Interested | 1 | 2 | 1 | 1 | 3 | 2 | 1 | 4 | 3 | 2 | 4 | 2 | 1 | 4 | 3 | 2 | 4 | 2 | 2 | 1 | -1 | 1 | 1 | 0 |
Excited | 2 | 3 | 1 | 2 | 2 | 0 | 1 | 4 | 3 | 2 | 4 | 2 | 2 | 4 | 2 | 2 | 4 | 2 | 2 | 1 | -1 | 2 | 1 | -1 |
Strong | 2 | 4 | 2 | 2 | 3 | 1 | 1 | 4 | 3 | 3 | 4 | 1 | 1 | 5 | 4 | 2 | 5 | 3 | 2 | 1 | -1 | 2 | 2 | 0 |
Enthusiastic | 1 | 2 | 1 | 2 | 3 | 1 | 3 | 3 | 0 | 2 | 4 | 2 | 1 | 3 | 2 | 2 | 4 | 2 | 2 | 2 | 0 | 2 | 2 | 0 |
Proud | 2 | 3 | 1 | 2 | 2 | 0 | 2 | 4 | 2 | 2 | 5 | 3 | 2 | 4 | 2 | 2 | 5 | 3 | 1 | 1 | 0 | 1 | 1 | 0 |
Alert | 3 | 3 | 0 | 1 | 2 | 1 | 2 | 4 | 2 | 2 | 4 | 2 | 2 | 3 | 1 | 1 | 3 | 2 | 5 | 5 | 0 | 2 | 2 | 0 |
Inspired | 2 | 3 | 1 | 2 | 1 | -1 | 1 | 4 | 3 | 2 | 5 | 3 | 1 | 4 | 3 | 2 | 5 | 3 | 1 | 1 | 0 | 2 | 2 | 0 |
Determined | 2 | 3 | 1 | 2 | 1 | -1 | 1 | 4 | 3 | 2 | 4 | 2 | 1 | 4 | 3 | 2 | 5 | 3 | 1 | 1 | 0 | 2 | 2 | 0 |
Attentive | 3 | 4 | 1 | 2 | 3 | 1 | 1 | 4 | 3 | 2 | 4 | 2 | 2 | 5 | 3 | 2 | 4 | 2 | 1 | 2 | 1 | 2 | 2 | 0 |
Active | 2 | 3 | 1 | 2 | 2 | 0 | 1 | 4 | 3 | 3 | 4 | 1 | 2 | 4 | 2 | 2 | 4 | 2 | 1 | 1 | 0 | 2 | 1 | -1 |
Distressed | 3 | 4 | 1 | 4 | 2 | -2 | 4 | 2 | -2 | 4 | 1 | -3 | 4 | 2 | -2 | 4 | 2 | -2 | 4 | 4 | 0 | 4 | 4 | 0 |
Upset | 4 | 4 | 0 | 5 | 4 | -1 | 4 | 2 | -2 | 4 | 2 | -2 | 4 | 1 | -3 | 4 | 1 | -3 | 5 | 5 | 0 | 4 | 4 | 0 |
Guilty | 3 | 2 | -1 | 4 | 4 | 0 | 4 | 2 | -2 | 3 | 2 | -1 | 3 | 1 | -2 | 3 | 1 | -2 | 3 | 3 | 0 | 3 | 3 | 0 |
Scared | 4 | 4 | 0 | 4 | 3 | -1 | 5 | 1 | -4 | 4 | 2 | -2 | 4 | 1 | -3 | 4 | 1 | -3 | 3 | 3 | 0 | 3 | 3 | 0 |
Hostile | 3 | 2 | -1 | 1 | 2 | 1 | 5 | 2 | -3 | 4 | 2 | -2 | 4 | 1 | -3 | 3 | 1 | -2 | 3 | 4 | 1 | 4 | 4 | 0 |
Irritable | 3 | 3 | 0 | 1 | 3 | 2 | 5 | 2 | -3 | 4 | 1 | -3 | 4 | 2 | -2 | 4 | 1 | -3 | 4 | 3 | -1 | 3 | 4 | 1 |
Ashamed | 4 | 3 | -1 | 5 | 4 | -1 | 4 | 3 | -1 | 3 | 3 | 0 | 4 | 2 | -2 | 3 | 2 | -1 | 3 | 3 | 0 | 3 | 2 | -1 |
Nervous | 4 | 3 | -1 | 4 | 4 | 0 | 4 | 2 | -2 | 4 | 2 | -2 | 4 | 1 | -3 | 4 | 1 | -3 | 4 | 5 | 1 | 3 | 2 | -1 |
Jittery | 4 | 3 | -1 | 4 | 3 | -1 | 4 | 2 | -2 | 4 | 2 | -2 | 4 | 1 | -3 | 4 | 1 | -3 | 4 | 5 | 1 | 4 | 5 | 1 |
Afraid | 3 | 2 | -1 | 4 | 4 | 0 | 4 | 2 | -2 | 3 | 1 | -2 | 5 | 2 | -3 | 4 | 2 | -2 | 3 | 4 | 1 | 3 | 2 | -1 |
Appendix D Chinese Experimental Results
In this section, we present the results of the experiments conducted in Chinese, with the same experimental setup except for the language difference.
Table 8 illustrates the results of the scene ablation experiment conducted in Chinese, indicating that our system exhibits stability across different scenarios. Figure 6 presents the results of the agent ablation experiment, indicating that the absence of the agent leads to a significant decline in outcomes regardless of the language used. This further demonstrates the rationality of our architecture.
@@ Appendix D Chinese Experimental Results
Appendix E Evaluation Metrics Description
In this section, we describe the six evaluation metrics of the main experiment in detail from the aspects of basic description, targeted dimensions, and references.
Metric | Description | Dimension | Reference |
Coherence | Assesses if the generated content is logical and transitions smoothly. | Focus on the quality of content generation: Evaluate the continuity of model-generated content, including plot coherence, logical consistency, and contextual flow between preceding and following texts. | (Kumaran et al., 2023) |
Immersion | Measures whether the user feels fully engaged and captivated by the interaction. | Focus on game scenario construction: Assess the level of player immersion within the game narrative. | (Jennett et al., 2008) |
Engagement | Evaluates if the system encourages sustained and meaningful interaction. | Focus on game interaction: Measure the degree of interactivity between the player and the game. | (Zhang et al., 2023) |
Emotional Relief | Measures if the interaction reduces user stress or anxiety. | Focus on the effectiveness of the framework: Determine whether the framework successfully alleviates the user’s emotions. | (Gross;, 1998) |
Satisfaction | Reflects the user’s overall contentment with the system. | Focus on user experience: Assess overall user satisfaction with the system. | (Nacke and Drachen, 2011) |
Interest | Assesses whether the content grabs attention and sparks curiosity. | Focus on user experience: Indicate the appeal of the generated interactive fiction games | (Nacke and Drachen, 2011) |
Appendix F Prompt Templates